I need help with a word problem on my assignment:
A nut wholesaler sells cashews for $3.50 a pound, and peanuts for $2 a pound. If he was going to create a 60 pound mixture to sell for $3 a pound, how many cups of each type of nuts would he need?

Answer by richwmiller(17219)

3.5c+2p=60*3
c+p=60
c=40
p=20